OpenStudy (anonymous):

What will the exponent be when you simplify the following problem? b^6 x b^3 = b^?

OpenStudy (anonymous):

add 'em

OpenStudy (2bornot2b):

9

OpenStudy (anonymous):

b^9

OpenStudy (anonymous):

First law of indices: \[a ^{m} \times a ^{n} = a ^{m \times n}\]

OpenStudy (anonymous):

No similarilli, it's a^(m+n)

OpenStudy (anonymous):

Ooops, I mean \[a ^{m+n}\]

OpenStudy (anonymous):

That's better ;) And sorry I typed your name completely wrong haha.

OpenStudy (anonymous):

The first law is. \[a ^{m} \times a^{n}=a ^{m+n}\]